### v1

# Stochastic Kinetics of Transcription: Analysis and Computation

Abstract
Stochastic gene expression is a fundamental yet complicated problem in biophysics, aiming to quantitatively describe the central dogma. Lack of thorough analysis and computationally efficient approach for fine-grained transcription models has confined relevant research to canonical Telegraph model. We study a general model of transcription and exactly solve the time-dependent chemical master equation. An approach highlighting computational efficiency is presented and analyzed for stationary distribution of mRNA molecules. Our result features generality, outstanding computational performance and explicit error estimation, unifying existing theoretical results on transcription kinetics and exhibiting applicability in practical implementation.
1 Introduction
The central dogma states that gene expression consists of transcription and translation. Transcription, synthesis of messenger RNA (mRNA) directed by genetic information encoded in DNA, is fundamental to almost all biophysical processes. Inherent stochasticity of transcription stems from stochastic nature of biochemical reactions, and becomes dominant due to low copy number of mRNA and existence of DNA as individual molecules [1]. This intrinsic fluctuation contributes prominently to the phenotypic heterogeneity [2], and is studied extensively both experimentally and mathematically [3].
The canonical mathematical formulation of transcription is Telegraph model, where gene switches memorylessly between active and inactive states [4]. Telegraph model is widely used to quantitatively study the bursting kinetics [3, 5], because it admits a concise and computationally tractable expression of steady-state distribution of mRNA molecules. However, Telegraph model is an overly simplified description of transcription [6]. Substantial experimental evidence suggests that transcription is a complex process governed by intricate mechanism [7]. In particular, holding time of either state is observed to be nonexponential [5], implying the complex structure of gene states. Numerous refined mathematical models of transcription are proposed and analyzed [8, 9, 10, 11, 12], but few obtains comprehensive and pragmatic results. More precisely, most existing studies are confined to models with specific gene-state structure and computationally feasible results are restricted to low-order summary statistics.
Essentially, the difficulty lies in the intractability of chemical master equation (CME) [13]. In general, CME can rarely anticipate a time-dependent exact solution and is usually studied approximately [14] or numerically [15, 16]. Among the methods, stochastic simulation algorithm (also known as Gillespie algorithm or SSA) [15] is the only exact approach. A variety of improvements of the original SSA have been devised [15]. However, in general, such stochastic simulation methods are computationally expensive and rely heavily on empirical exploration, thus may become inefficient in practical implementation. Approaches based on queueing theory are also exploited to provide theoretical results for an extensive range of stochastic gene expression problems, circumventing chemical master equation [17, 18]. But shortcomings include lack of physical interpretation and high computational cost. Even if CME is analyzed directly, we usually consider only the steady-state solution, setting the time derivatives to zero. In such case, the existence of equilibrium is presumed, which is not necessarily valid. Even if the steady-state solution is obtained, we have limited information about long time behavior of reaction kinetics since whether the system approaches the equilibrium and the rate of approaching are obscure.
In this Letter, we analyze a general model of transcription with complicated gene-state architecture and state transition upon transcription. We constructively present the exact time-dependent solution to the corresponding CME via power series method for solving differential equation. It is then shown that the probability distribution of mRNA molecules can be efficiently computed at steady state based on binomial moments. Explicit error estimation for practical implementation are also provided. Our approach is generally based on CME and generating function method, involving original matrix calculus techniques. Our result features generality, outstanding computational performance and explicit error estimation.
This Letter is structured as follows. We first present the main results as stated above. The detailed derivations can be found in the Supplementary Material. Then we derive two special cases from our general result: canonical Telegraph model [4] and the circumstance under renewal condition [17]. With mathematical reduction, we show the consistency of our result with previous studies. Verification of our analytical result by SSA is also performed. Finally, we briefly show how intrinsic and extrinsic noise can be collaboratively studied in our model and generalize the Heavy-Tailed Law proposed in [19].
2 Main Results
In this Letter, we mainly study a fine-grained model of transcription with complicated gene-state architecture, as illustrated in the following reaction scheme (1).
[…]
$\mathcal{S}_{i}(1\leq i\leq N)$ denotes $N$ gene-states; M denotes mRNA molecule. Reactions in the first row represent transitions among different states without transcription, while reactions in the second row are transitions with production of single mRNA molecule. The reaction in the last denotes hydrolysis of mRNA. In this model, arbitrarily many states of gene are allowed and transitions occur with or without production of mRNA molecules. Considering the standard presumption for single-step biochemical reaction [20], all the transitions are supposed to be Markovian. Let $a_{i,j}(1\leq i\neq j\leq N)$ denotes the transition rate without transcription from state $i$ to $j$ and $b_{i,j}(1\leq i,j\leq N)$ the rate with transcription from state $i$ to $j$. Hydrolysis of mRNA molecules is assumed to be a first-order Poisson process independent of gene-state, with parameter $\delta$. Let $\mathbb{P}_{(i,j)}(m,t)$ denote the conditional probability that $m$ mRNA molecules are present in the system and gene happens to be in state $j$ at time $t$, assuming gene-state to be $i$ at time 00. The CME corresponding to (1) is ($\forall$ $1\leq i\leq N$ and $m\in\mathbb{N}$)
$\frac{\partial}{\partial t}\mathbb{P}_{(i,j)}(m,t)=$ $\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{N}a_{k,j}\mathbb{P}_{(i,k)}(m,t)+\sum_{k=1}^{N}b_{k,j}\mathbb{P}_{(i,k)}(m-1,t)$ (2)
$+(m+1)\delta\mathbb{P}_{(i,j)}(m+1,t)-m\delta\mathbb{P}_{(i,j)}(m,t)$
$-\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{N}a_{j,k}\mathbb{P}_{(i,j)}(m,t)-\sum_{k=1}^{N}b_{j,k}\mathbb{P}_{(i,j)}(m,t).$
$\mathbb{P}_{(i,j)}(-1,t)$ is conventionally taken as null. Let $a_{i,i}:=-\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N}a_{i,k}-\sum_{k=1}^{N}b_{i,k}$ and $D_{0}:=(a_{i,j})_{N\times N},D_{1}:=(b_{i,j})_{N\times N}$ be two parameter matrices. Define the matrix-form time-dependent generating function to be $\mathcal{G}(t,z)$, whose $(i,j)$-th element is the generating function of $\{\mathbb{P}_{(i,j)}(m,t)\}_{m\in\mathbb{N}}$, namely,
$\left[\mathcal{G}(t,z)\right]_{(i,j)}:=\sum_{m=0}^{\infty}z^{m}\mathbb{P}_{(i,j)}(m,t).$
By rearranging (2), we have partial differential equation
$\frac{\partial}{\partial t}\mathcal{G}(t,z)=\mathcal{G}(t,z)D_{0}+z\mathcal{G}(t,z)D_{1}+(1-z)\delta\frac{\partial}{\partial z}\mathcal{G}(t,z).$ (3)
Without loss of generality, assume no mRNA molecule is present in the system at time 00. Inspired by power series method for solving differential equation, we now constructively present the analytical expression of $\mathcal{G}(t,z)$, the time-dependent solution to (3) under initial condition $\mathcal{G}(0,z)=\bm{I}_{N}$.
$\mathcal{G}(t,z)=\mathrm{e}^{D_{0}t}+\sum_{k=1}^{\infty}\int_{\Omega_{k}}$ $\prod_{i=1}^{k}\left[1+(z-1)\mathrm{e}^{-\delta(t-t_{i})}\right]$ (4)
$\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}d\bm{t},$
where $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots t_{k}\leq t\}\subseteq\mathbb{R}^{k}$. A detailed proof is in the Supplementary Material [21]. Similar results can be found within the field of queueing theory [22]. We note that our derivation is based on CME and differs essentially from those in queueing theory. The existence of an analytical expression of $\mathcal{G}(t,z)$ reveals why relatively elegant results can be expected by analyzing specific models through different approaches [8, 9, 10]. Though theoretically satisfactory, (4) is too complicated for practical implementation. For mathematical and computational tractability, we instead consider the binomial moment of mRNA molecules in the time limit. Note that $D:=D_{0}+D_{1}$ is a $Q$-matrix, meaning off-diagonal elements are nonnegative and row sum is all null. Assume $D$ to be irreducible throughout the Letter. For simplicity, we assume the underlying Markov chain of gene-state has converged to equilibrium, and take the invariant distribution to be $\bm{\pi}$. This assumption is reasonable if the architecture of gene-state reconstruct substantially slow compared with transcription kinetics. We now cease tracking microstate information of gene and focus on the copy number of mRNA. By summing over current states and setting the initial distribution to be $\bm{\pi}$, we define scalar binomial moments and probability distribution of mRNA molecules ($m,n\in\mathbb{N}$)
$B_{m}:=\frac{1}{m!}\lim_{t\rightarrow\infty}\left[\frac{\partial^{m}}{\partial z^{m}}\bm{\pi}^{T}\mathcal{G}(t,z)\bm{1}\bigg{|}_{z=1}\right]\,\,\text{and}\,\,\mathcal{P}_{n}:=\frac{1}{n!}\lim_{t\rightarrow\infty}\left[\frac{\partial^{n}}{\partial z^{n}}\bm{\pi}^{T}\mathcal{G}(t,z)\bm{1}\bigg{|}_{z=0}\right]$
should they exist, where $\bm{1}$ is a $N\times 1$ column vector of all ones. We assert that stationary binomial moments $\{B_{m}\}_{m\geq 1}$ not only exist but have the following compact expression
$B_{m}$ $=\frac{1}{m\delta}\bm{\pi}^{T}D_{1}\prod_{k=1}^{m-1}\left[\left(k\delta\bm{I}_{N}-D\right)^{-1}D_{1}\right]\bm{1},\quad m\geq 1,$ (5)
where $\bm{I}_{N}$ is the identity matrix. Note $\lambda\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for $\lambda>0$, thus nonsingular by Lévy-Desplanques theorem [23]. The derivation of (5) is in the Supplementary Material [21]. The special case where $D_{1}$ is diagonal has already been studied as a queueing system $MMPP/M/\infty$ [17], and (5) agrees with existing results [24]. As demonstrated in [25], binomial moments can be more tractable than probability distribution itself in the study of reaction system.
Furthermore, we can estimate the convergence of $\{B_{m}\}_{m\in\mathbb{N}}$. Note that $B_{0}$ is always unity by definition. For a strictly row diagonally dominant matrix $C=(c_{i,j})_{N\times N}$ with $\Theta$ defined as $\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$, $\lVert C^{-1}\rVert_{\infty}$ is bounded by $\Theta^{-1}$ from above (an analogue to Theorem 4.1.2 in [26]). The following result can confirm convergence and further delineate the velocity of convergence. To specify,
$B_{m}\leq\frac{1}{m!}\left(\frac{\lVert D_{1}\rVert_{\infty}}{\delta}\right)^{m}.$ (6)
In particular, $\{B_{m}\}_{m\in\mathbb{N}}$ converges to zero.
Exact probability distribution of mRNA molecules $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is uniquely determined by binomial moments $\{B_{m}\}_{m\in\mathbb{N}}$ and can be given by the identity
$\mathcal{P}_{n}=\sum^{\infty}_{m=n}\left(-1\right)^{m-n}\binom{m}{n}B_{m},\quad n\in\mathbb{N},$ (7)
where $\binom{m}{n}$ is the combinatorial number. Substituting (5) into (7), a theoretically exact and computationally feasible approach to obtaining stationary distribution of mRNA copy number is established. Additionally, we demonstrate that series (7) converges, proving the existence of a stationary distribution of mRNA population given arbitrary parameters in (1). The convergence of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ towards zero is also guaranteed. To quantify the asymptotic behavior of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$, we have
$\mathcal{P}_{n}\leq\frac{1}{n!}\mathrm{e}^{\lVert D_{1}\rVert_{\infty}/\delta}\left(\frac{\lVert D_{1}\rVert_{\infty}}{\delta}\right)^{n},\quad n\in\mathbb{N}.$ (8)
Since the upper bound converges to zero as $n$ grows, sequence $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ tends to null. Note also that this bound becomes practical only when $n$ is relatively large. As a series, evaluating the value of $\mathcal{P}_{n}$ based on (7) involves determining the truncation of terms. To ensure satisfactory accuracy, a priori estimation of truncation error is expected and is substantially favored over empirical choice. Nonetheless, based on our knowledge, none of existing approaches to analyzing stochastic gene expression model effectively achieve this objective. In contrast, in our model, we are able to give an explicit upper bound for truncation error. Let $\widehat{\mathcal{P}_{n}}$ denote the computed value given the truncation point $M$. Assume $M>\max\{2n+2,8\mathrm{e}\left(\lVert D_{1}\rVert_{\infty}/\delta\right)^{2}\}$. Then
$\mid\mathcal{P}_{n}-\widehat{\mathcal{P}_{n}}\mid\leq\frac{1}{\sqrt{2\pi\mathrm{e}}n!2^{M-1}}.$ (9)
Conventionally, when analyzing biochemical reaction systems like (1), we assume in advance the existence of equilibrium, and calculate the steady-state solution with all the time derivatives in the CME set to zero. In this Letter, instead, we present directly an analytical formulation of the time-dependent solution and then take the limit with respect to time. Thereby, our results additionally illustrates that the reaction system (1) possesses a unique steady state and always converges to equilibrium after a sufficiently long period of time. The rate of approaching the equilibrium can also be obtained by studying the time-dependent solution (4). Moreover, explicit upper bounds (6)(8)(9) can be exploited to design a predictable computational framework for calculating the summary statistics of arbitrarily high order or the probability distribution of mRNA copy number.
Figure 1: Upper Bound of Binomial Moments and Probability Distribution of mRNA molecules: In the left illustration, binomial moments and upper bound are computed according to (5) and (6), respectively. In the right illustration, the upper bound of probability distribution is given by (8). Without loss of generality, we normalize all the parameters by dividing $\delta$. Hence $\delta$ is set as unity in the following examples. In the above illustrations, parameters are both set as $D_{0}=$ $\begin{pmatrix}-7&2&3&0\\ 3&-12&1&5\\ 1&2&-12&6\\ 8&8&0&-19\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&1&0&0\\ 1&0&1&1\\ 0&1&1&1\\ 1&1&0&1\end{pmatrix}$
3 Verification by Simulation and Connection with some Known Results
3.1 Stochastic Simulation
Stochastic simulation algorithm serves as a universal and exact approach to studying chemical reactions. We first verify our analytical result (5)(7) with stochastic simulation [27] of the reaction system (1). As Figure 2 demonstrates, our theoretical result of steady-state distribution of mRNA molecules agrees well with output of simulations.
Figure 2: Probability Distribution of mRNA Molecules Obtained by Stochastic Simulation and Our Theoretical Result: To perform SSA, we implement Python package GillesPy2 [27]. In Example $1$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-16.01\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&15\end{pmatrix}$ The histogram is generated based on $50000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results presented in the Letter. In Example $2$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.2&0.1&0&0&0.1\\ 0.1&-2.7&0.5&0&0.1\\ 0.1&0.2&-5.4&0&0.1\\ 0.1&0.1&0.3&-3.5&0\\ 0.1&0.1&0.1&0.1&-15.4\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}0&0&1&0&1\\ 1&1&0&0&0\\ 1&2&1&1&0\\ 1&0&0&1&1\\ 0&0&0&0&15\end{pmatrix}$ The histogram is generated based on $100000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results. Each trajectory is truncated at dimensionless time $t=100$.
Additionally, our model (1) possesses a significant degree of generality, covering many previous results. In this Letter, we focus on only two exemplars: the classic Telegraph model [4] and the special case where transcription becomes a renewal process [17].
3.2 Special Case I: Telegraph model
Taking $D_{0}=\begin{pmatrix}-k_{1}-\lambda&k_{1}\\ k_{2}&-k_{2}\end{pmatrix}$ and $D_{1}=\text{diag}[\lambda,0]$, our model (1) reduces to the classic Telegraph model [4], the pioneering work in the field of stochastic gene expression. It is clear that $\bm{\pi}=\left[k_{2}/(k_{1}+k_{2}),k_{1}/(k_{1}+k_{2})\right]^{T}$ and
$\left(k\delta\bm{I}_{2}-D\right)^{-1}=\frac{1}{k\delta(k\delta+k_{1}+k_{2})}\begin{pmatrix}k_{2}+k\delta&-k_{1}\\ -k_{2}&k_{1}+k\delta\end{pmatrix}.$ (10)
Substituting into (5), we have
$B_{m}$ $=\frac{1}{m!}\left(\frac{\lambda}{\delta}\right)^{m}\frac{(k_{2}/\delta)^{(m)}}{(k_{1}/\delta+k_{2}/\delta)^{(m)}},$ (11)
where $x^{(m)}:=\Gamma(x+m)/\Gamma(x)$ is the Pochhammer symbol. $\Gamma(x)$ is Gamma function. By substituting (11) into (7), it can be derived that
$\mathcal{P}_{n}$ $=\frac{1}{n!}\left(\frac{\lambda}{\delta}\right)^{n}\frac{(k_{2}/\delta)^{(n)}}{(k_{1}/\delta+k_{2}/\delta)^{(n)}}\quad\mathllap{{}_{1}}F_{1}\left(\frac{k_{2}}{\delta}+n,\frac{k_{1}+k_{2}}{\delta}+n;-\frac{\lambda}{\delta}\right),$ (12)
where  $\mathllap{{}_{1}}F_{1}(a,b;z):=\sum_{n=0}^{\infty}\frac{a^{(n)}}{b^{(n)}}\frac{z^{n}}{n!}$ denotes the confluent hypergeometric function. (12) is consistent with previous result [4, 8]. A detailed calculation can be found in the Supplementary Material [21].
Although formulation of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ in (12) seems formally compact, the confluent hypergeometric function is, in essence, an abbreviation of an infinite series. Evaluating the value of confluent hypergeometric function at given point possesses approximately the same computational complexity as one calculation in (7), and it involves truncation error as well. From this standpoint, (12) holds limited advantage, both mathematically and computationally, over reconstructing $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ from binomial moments according to (7). Likewise, this observation holds for other known results involving generalized hypergeometric function [8, 9, 10, 17].
3.3 Special Case II: Under Renewal Condition
Previous studies predominantly focus on the case where transcription emerges as a renewal process and queueing theory is introduced to obtain powerful results in a relatively general framework [18]. Transcription in (1) becomes a renewal process if and only if the following renewal condition [17] holds:
$\text{if}\;\;\sum_{k=1}^{N}b_{i,k}\neq 0,\;\;\text{then}\;\;\frac{b_{i,k}}{\sum_{k=1}^{N}b_{i,k}}\;\;\text{is independent of}\;\;i.$ (13)
Assuming $D_{1}\neq\bm{0}_{N\times N}$, the renewal condition (13) is equivalent to $\text{rank}\,D_{1}=1$. Under the renewal condition, our model (1) can be interpreted otherwise as a queueing system $GI/M/\infty$ with probability density function of inter-arrival time [17]
$f(t)=\bm{\kappa}^{T}\mathrm{e}^{D_{0}t}D_{1}\bm{1},$ (14)
where $\bm{\kappa}=(\kappa_{j})^{T}=(\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}})^{T}$ as long as $\sum_{j=1}^{N}b_{i,j}\neq 0$.
Queueing system with arrivals following a renewal process has been thoroughly studied [28]. The tractable nature of renewal process yields satisfactory theoretical results. For queueing system $GI/M/\infty$, the stationary binomial moments can be given by [28]
$B_{m}=\frac{1}{\alpha m\delta}\prod_{i=1}^{m-1}\frac{\mathcal{L}[f](i\delta)}{1-\mathcal{L}[f](i\delta)},\quad m\geq 1,$ (15)
[…]
Assuming nonsingularity of $D_{0}$, binomial moments given in (15) can be reduced to our result given in (5). The details can be found in the Supplementary Material [21]. We note that, under the assumption $D$ is irreducible and $\text{rank}\,D_{1}=1$, $D_{0}$ is nonsingular, which is equivalent to convergence of $\alpha$. A proof is given in the Supplementary Material [21], which can be seen as a generalization of Taussky theorem [23].
3.4 Computational Performance
Figure 3: Computation Time of Binomial Moments and SSA: Computation time is evaluated for increasing model complexity and different approaches. As dimension of $D_{0}$ or $D_{1}$ increases, number of reaction pathways escalates, making the model more complicated. In the above illustration, computation time pertains to an exemplar where $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$. In the left figure, three different approaches are employed to models with dimension $i\,(3\leq i\leq 8)$; in the right one, two approaches are employed to model with dimension $5j\,(1\leq j\leq 10)$. Curve of our result is computation time of the first $200$ binomial moments based on (5). Curve of renewal process is computation time of the first $200$ binomial moments based on (15). SSA is performed for $1000$ trajectories and truncated at dimensionless time $t=100$ (GillesPy2 [27] is implemented with Python solver). Time is evaluated for three replicate computations and averaged to mitigate randomness.
Our result for (1) is performed mainly in the framework of matrix computation, which is one of the most efficient and well-developed area in scientific computation. As we illustrate in Figure 3, our result stands out among several existing approaches in computational efficiency, where we compute the first $200$ binomial moments based on (5) and (15), respectively, and implement SSA. Separation of time scales can be distinctly observed with respect to our result, the renewal process approach and SSA. In particular, our result remains computationally feasible for high-dimensional model while other approaches certainly fail. Moreover, our result also admits explicit error estimation, which is, as far as we know, unattainable through other existing methods. Additionally, accuracy, stability and further improvement on efficiency can be pursued by exploiting the structure of matrix in (5), especially $k\delta\bm{I}_{N}-D$. For example, $k\delta\bm{I}_{N}-D$ is diagonally dominant, thus can perform LU factorization without pivoting (Theorem $4.1.1$ in [26]). It involves $\frac{8}{3}N^{3}$ flops to calculate $\left(k\delta\bm{I}_{N}-D\right)^{-1}$ for given $k$ and $\frac{20}{3}(m-1)N^{3}$ flops to calculate $B_{m}$ in general. Recursive structure in (5) can also be utilized to reduce overall complexity.
In comparison, SSA has two notable shortcomings. Firstly, SSA is computationally intensive and thus impractical to implement for high-dimensional reaction system. The second is that SSA relies heavily on empirical insights. Number of simulations and truncation time point to generate a stationary distribution are both empirically determined, hence the implementation requires repeated trials. The renewal process approach also has several shortcomings. Primarily, Laplace transform of inter-arrival time probability density distribution can be mathematically challenging to obtain given specific gene expression model. Inter-arrival time distribution also conceals physical interpretation of gene-state architecture while parameters in (1) have clear biophysical meanings. Additionally, results based on renewal process is computational infeasible excluding a few special cases. In general, computational complexity become exceedingly high if high-order binomial moment is involved.
4 Mathematical Description of Noise in Gene Expression
Noise in gene expression within a supposedly identical ensemble arises from multiple sources [2, 3]. Quantitative understanding of generation and propagation of noise is one of the primary objectives of studying stochastic gene expression, and noise in transcription can predominantly contribute to protein-level fluctuation [29]. Experimentally, noise can be categorized into intrinsic and extrinsic noise. Intrinsic noise is the fluctuation originating from inherent stochasticity of biochemical reactions. Deterministic description fails since number of molecules involved in gene expression can be relatively low. Our model (1) provides a general framework to investigate origins and consequences of intrinsic noise in gene expression. Based on our result (5), Fano factor, the ratio of variance to mean of mRNA copy number, is given by
$FF=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}}=\frac{\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{\bm{\pi}^{T}D_{1}\bm{1}}+1-\frac{\bm{\pi}^{T}D_{1}\bm{1}}{\delta}.$ (16)
Intrinsic noise, defined as the ratio of variance to mean square, is
$\eta^{2}=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}^{2}}=\frac{\delta\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{(\bm{\pi}^{T}D_{1}\bm{1})^{2}}+\frac{\delta}{\bm{\pi}^{T}D_{1}\bm{1}}-1.$ (17)
For specific gene-state structure where $(\delta\bm{I}_{N}-D)^{-1}$ has explicit expression, further reduction of Fano factor and intrinsic noise may be performed. Consistency of (16) and (17) with previous results can be verified [9, 17].
Extrinsic noise refers to fluctuation caused by environmental differences among cells, and can be further classified as local and global extrinsic noise [2]. Although intrinsic noise can easily be grounded in mathematical formulation, as our model (1) does, incorporating extrinsic noise into chemical mater equation is conventionally considered to be formidable. One insightful approach is to utilize the parametric nature of the model [19]. Parameters, which are $D_{0},D_{1}$ and $\delta$ in our case, are assumed to follow some probability distribution with density function $\rho(\bm{\theta})$, representing effects of extrinsic fluctuation. Since binomial moments have concise expression given by (5), posing the perturbation on binomial moments and then reconstruct probability distribution is more tractable than posing perturbation directly on probability distribution as in [19]. Thereby, the expected binomial moments of mRNA molecules subject to extrinsic noise can be formulated as
[…]
where $\bm{\theta}$ abbreviates the vector of $12N^{2}-N+$ 1 independent parameters in our model (1) and $\Omega$ denotes the parameter space $[0,\infty]^{2N^{2}-N+1}$. Substituting (18) into (7), perturbed probability distribution can be reconstructed. For specific $\rho(\bm{\theta})$ and simplified model, (18) may be reduced to closed form.
This framework can be used to apprehend the joint effects of intrinsic and extrinsic noise, which has been performed for some special cases of our model (1) [8, 19]. Heavy-Tailed Law [19] asserts that intrinsic noise alone cannot generate mRNA distributions that is heavy tailed, namely, the moment generating function [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)$ becomes infinity for $t>0$. Heavy-Tailed Law has been proposed intuitively for Leaky Telegraph model [19] and a generalized model [8]. In this Letter, we provide a mathematically rigorous proof of Heavy-Tailed Law in our model (1), generalizing previous results. More preciously, we prove the following inequality based on (8).
[$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)\leq\mathrm{e}^{2\lVert D_{1}\rVert_{\infty}/\delta}\times\mathcal{M}\left[\text{Pois}(\frac{\lVert D_{1}\rVert_{\infty}}{\delta})\right](t).$ (19)
For Poisson distribution, moment generating function is finite for any $t>0$, therefore so is $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ according to (19). Therefore $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is never heavy tailed. It follows that experimentally observed heavy-tailed distribution of mRNA copy number implies collaborative effect of intrinsic and extrinsic noise [19]. The mathematical proof of (19) is in Supplementary Material [21].
5 Discussion
Stochastic gene expression is a central problem in biophysics. Numerous mathematical theories have been developed to provide quantitative insights into specific models. In this Letter, we study a general kinematic model of transcription. The exact time-dependent solution to the chemical master equation is constructed. A computationally feasible approach based on binomial moments is presented and thoroughly analyzed. We examine two special cases and perform SSA to verify our results. Collaborative effect of intrinsic and extrinsic noise are discussed. A mathematically rigorous proof of generalized Heavy-Tailed Law is also provided. Compared with other existing methods, our approach features generality, exceptional computational efficiency as well as explicit error estimation. Results presented in this Letter boost further analysis of origin and variation of noise. Dependence of noise and Fano factor on different parameters should be further investigated. Note also that our model is theoretically motivated. When modeling specific type of gene, physical background should be exploited to constrain model complexity appropriately. Statistical inference can further be employed by fitting our model to experimental data, and finally reveal the underlying mechanisms of transcription.
Acknowledgments
Y. L. is supported by FDUROP (Fudan Undergraduate Research Opportunities Program). Y. Z. is supported by National Key R&\&&D Program of China (2024YFA1012401), the Science and Technology Commission of Shanghai Municipality (23JC1400501), and Natural Science Foundation of China (12241103).

### v2

@@
-Stochastic Kinetics of Transcription: Analysis and Computation
+Stochastic Kinetics of mRNA Molecules in a General Transcription Model
 Abstract
@@
 An approach highlighting computational efficiency is presented and analyzed for stationary distribution of mRNA molecules.
-Our result features generality, outstanding computational performance and explicit error estimation, unifying existing theoretical results on transcription kinetics and exhibiting applicability in practical implementation.
+Distribution of mRNA copy number is found to be upper-bounded by a constant multiple of Poisson distribution.
+Our result features generality, profundity and outstanding computational performance, unifying existing theoretical results on transcription kinetics and exhibiting broader applicability in practical implementation.
 1 Introduction
@@
 Transcription, synthesis of messenger RNA (mRNA) directed by genetic information encoded in DNA, is fundamental to almost all biophysical processes.
-Inherent stochasticity of transcription stems from stochastic nature of biochemical reactions, and becomes dominant due to low copy number of mRNA and existence of DNA as individual molecules [1].
+Inherent stochasticity of transcription stems from stochastic nature of biochemical reactions, and becomes dominant due to low copy number of mRNA and existence of DNA as individual molecules [11, 21].
 This intrinsic fluctuation contributes prominently to the phenotypic heterogeneity [32], and is studied extensively both experimentally and mathematically [3].
+In particular, an overarching theoretical perspective is provided in [29].
 The canonical mathematical formulation of transcription is Telegraph model, where gene switches memorylessly between active and inactive states [30].
-Telegraph model is widely used to quantitatively study the bursting kinetics [3, 5], because it admits a concise and computationally tractable expression of steady-state distribution of mRNA molecules.
+Telegraph model is widely used to quantitatively study the bursting kinetics [26, 36] and transcriptional regulation [31], because it admits a concise and computationally tractable expression of steady-state distribution of mRNA molecules.
 However, Telegraph model is an overly simplified description of transcription [18].
@@
 In particular, holding time of either state is observed to be nonexponential [36], implying the complex structure of gene states.
-Numerous refined mathematical models of transcription are proposed and analyzed [8, 9, 10, 11, 12], but few obtains comprehensive and pragmatic results.
+Numerous refined mathematical models of transcription are proposed and analyzed [4, 13, 14, 16, 20, 35, 38], but few obtains comprehensive and pragmatic results.
 More precisely, most existing studies are confined to models with specific gene-state structure and computationally feasible results are restricted to low-order summary statistics.
-Essentially, the difficulty lies in the intractability of chemical master equation (CME) [13].
-In general, CME can rarely anticipate a time-dependent exact solution and is usually studied approximately [14] or numerically [15, 16].
+Essentially, the difficulty lies in the intractability of chemical master equation (master equation or CME) [19].
+In general, CME can rarely anticipate a time-dependent exact solution [9] and is usually studied approximately [2, 33] or numerically [7, 12, 25].
 Among the methods, stochastic simulation algorithm (also known as Gillespie algorithm or SSA) [7] is the only exact approach.
@@
 Even if the steady-state solution is obtained, we have limited information about long time behavior of reaction kinetics since whether the system approaches the equilibrium and the rate of approaching are obscure.
-In this Letter, we analyze a general model of transcription with complicated gene-state architecture and state transition upon transcription.
+However, we demonstrate in this article that the chemical master equation of a general transcription model is exactly solvable.
+In this article, we analyze a general model of transcription with complicated gene-state architecture and state transition upon transcription.
 We constructively present the exact time-dependent solution to the corresponding CME via power series method for solving differential equation.
-It is then shown that the probability distribution of mRNA molecules can be efficiently computed at steady state based on binomial moments.
-Explicit error estimation for practical implementation are also provided.
+It is then shown that the probability density function of mRNA molecules can be effectively investigated at steady state based on binomial moments.
+Sharp upper bounds for binomial moments are given based on an analogue to Theorem 4.1.2 in [8].
+Subsequently, an upper bound for probability density function is also derived, which shows that probability density function of mRNA molecules is always bounded by a constant multiple of Poisson distribution.
+This observation is later used to generalize the Heavy-Tailed Law proposed in [10].
+Explicit error estimation for practical implementation is also provided.
 Our approach is generally based on CME and generating function method, involving original matrix calculus techniques.
-Our result features generality, outstanding computational performance and explicit error estimation.
-This Letter is structured as follows.
+Our result features generality, profundity and outstanding computational performance.
+This article is structured as follows.
 We first present the main results as stated above.
-The detailed derivations can be found in the Supplementary Material.
+The detailed derivations can be found in the Appendix.
 Then we derive two special cases from our general result: canonical Telegraph model [30] and the circumstance under renewal condition [37].
@@
 Verification of our analytical result by SSA is also performed.
-Finally, we briefly show how intrinsic and extrinsic noise can be collaboratively studied in our model and generalize the Heavy-Tailed Law proposed in [19].
+Finally, we briefly show how intrinsic and extrinsic noise can be collaboratively studied in our model and prove the generalized Heavy-Tailed Law.
 2 Main Results
-In this Letter, we mainly study a fine-grained model of transcription with complicated gene-state architecture, as illustrated in the following reaction scheme (1).
+In this article, we mainly study a fine-grained model of transcription with complicated gene-state architecture, as illustrated in the following reaction scheme (1).
 […]
@@
 In this model, arbitrarily many states of gene are allowed and transitions occur with or without production of mRNA molecules.
+If necessary, enlarge $N$ to secure that each transition represents one single-step biochemical reaction.
+Note that reactants not directly involved in the above process have been omitted to underscore the reaction structure.
+For instance, if translation of protein from mRNA and degradation of protein are considered [34], the marginal distribution of mRNA molecules will be exactly our result presented in this article.
 Considering the standard presumption for single-step biochemical reaction [6], all the transitions are supposed to be Markovian.
-Let $a_{i,j}(1\leq i\neq j\leq N)$ denotes the transition rate without transcription from state $i$ to $j$ and $b_{i,j}(1\leq i,j\leq N)$ the rate with transcription from state $i$ to $j$.
+Let $a_{i,j}(1\leq i\neq j\leq N)$ denote the transition rate without transcription from state $\mathcal{S}_{i}$ to $\mathcal{S}_{j}$ and $b_{i,j}(1\leq i,j\leq N)$ the rate with transcription from state $\mathcal{S}_{i}$ to $\mathcal{S}_{j}$.
 Hydrolysis of mRNA molecules is assumed to be a first-order Poisson process independent of gene-state, with parameter $\delta$.
-Let $\mathbb{P}_{(i,j)}(m,t)$ denote the conditional probability that $m$ mRNA molecules are present in the system and gene happens to be in state $j$ at time $t$, assuming gene-state to be $i$ at time 00.
-The CME corresponding to (1) is ($\forall$ $1\leq i\leq N$ and $m\in\mathbb{N}$)
+Let $\mathbb{P}_{(i,j)}(m,t)$ denote the conditional probability that $m$ mRNA molecules are present in the system and gene happens to be in state $\mathcal{S}_{j}$ at time $t$, assuming gene-state to be $\mathcal{S}_{i}$ at time 00.
+According to (1), number of mRNA molecules together with gene state is a continuous-time Markov chain with state space ×$\mathbb{N}\times\{1,2,\cdots,N\}$.
+The corresponding CME is ($\forall$ $1\leq i,j\leq N$ and $m\in\mathbb{N}$)
 $\frac{\partial}{\partial t}\mathbb{P}_{(i,j)}(m,t)=$ $\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{N}a_{k,j}\mathbb{P}_{(i,k)}(m,t)+\sum_{k=1}^{N}b_{k,j}\mathbb{P}_{(i,k)}(m-1,t)$ (2)
@@
 $\left[\mathcal{G}(t,z)\right]_{(i,j)}:=\sum_{m=0}^{\infty}z^{m}\mathbb{P}_{(i,j)}(m,t).$
-By rearranging (2), we have partial differential equation
+By rearranging (2), we have partial differential equation system
 $\frac{\partial}{\partial t}\mathcal{G}(t,z)=\mathcal{G}(t,z)D_{0}+z\mathcal{G}(t,z)D_{1}+(1-z)\delta\frac{\partial}{\partial z}\mathcal{G}(t,z).$ (3)
@@
 $\mathcal{G}(t,z)=\mathrm{e}^{D_{0}t}+\sum_{k=1}^{\infty}\int_{\Omega_{k}}$ $\prod_{i=1}^{k}\left[1+(z-1)\mathrm{e}^{-\delta(t-t_{i})}\right]$ (4)
-$\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}d\bm{t},$
-where $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots t_{k}\leq t\}\subseteq\mathbb{R}^{k}$.
-A detailed proof is in the Supplementary Material [21].
-Similar results can be found within the field of queueing theory [22].
+$\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}\mathrm{d}\bm{t},$
+where $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq t\}\subseteq\mathbb{R}^{k}$.
+A detailed proof is in Appendix A.
+Similar results can be found within the field of queueing theory [23], which provides physical interpretation of the expression (4).
 We note that our derivation is based on CME and differs essentially from those in queueing theory.
-The existence of an analytical expression of $\mathcal{G}(t,z)$ reveals why relatively elegant results can be expected by analyzing specific models through different approaches [8, 9, 10].
+The existence of an analytical expression of $\mathcal{G}(t,z)$ reveals why relatively elegant results about probability density function can be expected by analyzing specific models through different approaches [4, 13, 16, 38].
 Though theoretically satisfactory, (4) is too complicated for practical implementation.
 For mathematical and computational tractability, we instead consider the binomial moment of mRNA molecules in the time limit.
-Note that $D:=D_{0}+D_{1}$ is a $Q$-matrix, meaning off-diagonal elements are nonnegative and row sum is all null.
-Assume $D$ to be irreducible throughout the Letter.
+Note that $D:=D_{0}+D_{1}$ is a $Q$-matrix, meaning off-diagonal elements are nonnegative and row sum is null.
+Assume $D$ to be irreducible throughout the article.
 For simplicity, we assume the underlying Markov chain of gene-state has converged to equilibrium, and take the invariant distribution to be $\bm{\pi}$.
 This assumption is reasonable if the architecture of gene-state reconstruct substantially slow compared with transcription kinetics.
-We now cease tracking microstate information of gene and focus on the copy number of mRNA.
-By summing over current states and setting the initial distribution to be $\bm{\pi}$, we define scalar binomial moments and probability distribution of mRNA molecules ($m,n\in\mathbb{N}$)
+We now cease tracking microstates of gene and focus on the coarse-grained copy number of mRNA.
+By summing over current states and setting the initial distribution to be $\bm{\pi}$, we define scalar binomial moments and probability density function of mRNA molecules ($m,n\in\mathbb{N}$)
 $B_{m}:=\frac{1}{m!}\lim_{t\rightarrow\infty}\left[\frac{\partial^{m}}{\partial z^{m}}\bm{\pi}^{T}\mathcal{G}(t,z)\bm{1}\bigg{|}_{z=1}\right]\,\,\text{and}\,\,\mathcal{P}_{n}:=\frac{1}{n!}\lim_{t\rightarrow\infty}\left[\frac{\partial^{n}}{\partial z^{n}}\bm{\pi}^{T}\mathcal{G}(t,z)\bm{1}\bigg{|}_{z=0}\right]$
@@
 Note $\lambda\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for $\lambda>0$, thus nonsingular by Lévy-Desplanques theorem [15].
-The derivation of (5) is in the Supplementary Material [21].
+The derivation of (5) is in Appendix B.
 The special case where $D_{1}$ is diagonal has already been studied as a queueing system $MMPP/M/\infty$ [37], and (5) agrees with existing results [27].
-As demonstrated in [25], binomial moments can be more tractable than probability distribution itself in the study of reaction system.
-Furthermore, we can estimate the convergence of $\{B_{m}\}_{m\in\mathbb{N}}$.
+As demonstrated in [1], binomial moments can be more tractable than probability density function itself in the study of reaction system.
+Furthermore, we can estimate the long-time behavior of $\{B_{m}\}_{m\in\mathbb{N}}$.
 Note that $B_{0}$ is always unity by definition.
-For a strictly row diagonally dominant matrix $C=(c_{i,j})_{N\times N}$ with $\Theta$ defined as $\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$, $\lVert C^{-1}\rVert_{\infty}$ is bounded by $\Theta^{-1}$ from above (an analogue to Theorem 4.1.2 in [26]).
-The following result can confirm convergence and further delineate the velocity of convergence.
+For a strictly row diagonally dominant matrix $C=(c_{i,j})_{N\times N}$ with $\Theta$ defined as $\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$, $\lVert C^{-1}\rVert_{\infty}$ is bounded by $\Theta^{-1}$ from above (an analogue to Theorem 4.1.2 in Ref.
+[8]).
+The following result confirms convergence and further delineates the velocity of convergence.
 To specify,
@@
 In particular, $\{B_{m}\}_{m\in\mathbb{N}}$ converges to zero.
-Exact probability distribution of mRNA molecules $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is uniquely determined by binomial moments $\{B_{m}\}_{m\in\mathbb{N}}$ and can be given by the identity
+See the left panel in Figure 1.
+Proof of (6) is in Appendix C.
+Exact probability density function of mRNA molecules $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is uniquely determined by binomial moments $\{B_{m}\}_{m\in\mathbb{N}}$ and can be given according to the identity
 $\mathcal{P}_{n}=\sum^{\infty}_{m=n}\left(-1\right)^{m-n}\binom{m}{n}B_{m},\quad n\in\mathbb{N},$ (7)
@@
 Additionally, we demonstrate that series (7) converges, proving the existence of a stationary distribution of mRNA population given arbitrary parameters in (1).
-The convergence of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ towards zero is also guaranteed.
 To quantify the asymptotic behavior of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$, we have
 $\mathcal{P}_{n}\leq\frac{1}{n!}\mathrm{e}^{\lVert D_{1}\rVert_{\infty}/\delta}\left(\frac{\lVert D_{1}\rVert_{\infty}}{\delta}\right)^{n},\quad n\in\mathbb{N}.$ (8)
+This inequality demonstrates that probability density function of mRNA copy number is always bounded from above by a constant multiple of the density function of one Poisson distribution.
+To emphasize the usefulness of this upper bound, we utilize (8) to prove generalized Heavy-Tailed Law in section 4.
 Since the upper bound converges to zero as $n$ grows, sequence $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ tends to null.
 Note also that this bound becomes practical only when $n$ is relatively large.
+See the right illustration in Figure 1.
+Proof of (8) is in Appendix D.
 As a series, evaluating the value of $\mathcal{P}_{n}$ based on (7) involves determining the truncation of terms.
-To ensure satisfactory accuracy, a priori estimation of truncation error is expected and is substantially favored over empirical choice.
+To obtain satisfactory accuracy, a priori estimation of truncation error is expected and is substantially favored over empirical choice.
 Nonetheless, based on our knowledge, none of existing approaches to analyzing stochastic gene expression model effectively achieve this objective.
@@
 Then
-$\mid\mathcal{P}_{n}-\widehat{\mathcal{P}_{n}}\mid\leq\frac{1}{\sqrt{2\pi\mathrm{e}}n!2^{M-1}}.$ (9)
+$\mid\mathcal{P}_{n}-\widehat{\mathcal{P}_{n}}\mid\leq\frac{1}{\sqrt{2\pi\mathrm{e}}n!2^{M-1}},$ (9)
+which is proved in Appendix E.
+However, we note that, apart from truncation errors, floating-point errors can become extremely severe when reconstructing probability density function based on (7).
+Essentially the same issue has been observed and analyzed in [13].
 Conventionally, when analyzing biochemical reaction systems like (1), we assume in advance the existence of equilibrium, and calculate the steady-state solution with all the time derivatives in the CME set to zero.
-In this Letter, instead, we present directly an analytical formulation of the time-dependent solution and then take the limit with respect to time.
+In this article, instead, we construct directly an analytical formulation of the time-dependent solution and then take the limit with respect to time.
 Thereby, our results additionally illustrates that the reaction system (1) possesses a unique steady state and always converges to equilibrium after a sufficiently long period of time.
-The rate of approaching the equilibrium can also be obtained by studying the time-dependent solution (4).
-Moreover, explicit upper bounds (6)(8)(9) can be exploited to design a predictable computational framework for calculating the summary statistics of arbitrarily high order or the probability distribution of mRNA copy number.
-Figure 1: Upper Bound of Binomial Moments and Probability Distribution of mRNA molecules: In the left illustration, binomial moments and upper bound are computed according to (5) and (6), respectively.
-In the right illustration, the upper bound of probability distribution is given by (8).
+The rate of approaching the equilibrium can also be obtained by estimating the time-dependent solution (4).
+Moreover, explicit upper bounds (6)(8)(9) can be exploited to design a predictable computational framework for calculating the summary statistics of arbitrarily high order or the probability density function of mRNA copy number.
+Figure 1: Upper Bound of Binomial Moments and Probability Distribution of mRNA molecules: In the left panel, binomial moments and upper bound are computed according to (5) and (6), respectively.
+In the right panel, the upper bound of probability density function is given by (8).
 Without loss of generality, we normalize all the parameters by dividing $\delta$.
-Hence $\delta$ is set as unity in the following examples.
+Hence $\delta$ is set as unity in this and following examples.
 In the above illustrations, parameters are both set as $D_{0}=$ $\begin{pmatrix}-7&2&3&0\\ 3&-12&1&5\\ 1&2&-12&6\\ 8&8&0&-19\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&1&0&0\\ 1&0&1&1\\ 0&1&1&1\\ 1&1&0&1\end{pmatrix}$
-3 Verification by Simulation and Connection with some Known Results
-3.1 Stochastic Simulation
+3 Stochastic Simulation and Consistency with Known Results
 Stochastic simulation algorithm serves as a universal and exact approach to studying chemical reactions.
-We first verify our analytical result (5)(7) with stochastic simulation [27] of the reaction system (1).
+We first verify our analytical result (5)(7) with stochastic simulation of the reaction system (1).
 As Figure 2 demonstrates, our theoretical result of steady-state distribution of mRNA molecules agrees well with output of simulations.
-Figure 2: Probability Distribution of mRNA Molecules Obtained by Stochastic Simulation and Our Theoretical Result: To perform SSA, we implement Python package GillesPy2 [27].
-In Example $1$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-16.01\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&15\end{pmatrix}$ The histogram is generated based on $50000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results presented in the Letter.
+Figure 2: Probability Distribution of mRNA Molecules Obtained through Stochastic Simulation and Our Theoretical Result: To perform SSA, we implement Python package GillesPy2 [24].
+In Example $1$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-16.01\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&15\end{pmatrix}$ The histogram is generated based on $50000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results presented in the article.
 In Example $2$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.2&0.1&0&0&0.1\\ 0.1&-2.7&0.5&0&0.1\\ 0.1&0.2&-5.4&0&0.1\\ 0.1&0.1&0.3&-3.5&0\\ 0.1&0.1&0.1&0.1&-15.4\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}0&0&1&0&1\\ 1&1&0&0&0\\ 1&2&1&1&0\\ 1&0&0&1&1\\ 0&0&0&0&15\end{pmatrix}$ The histogram is generated based on $100000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results.
@@
 Additionally, our model (1) possesses a significant degree of generality, covering many previous results.
-In this Letter, we focus on only two exemplars: the classic Telegraph model [4] and the special case where transcription becomes a renewal process [17].
+In this article, we focus on only two exemplars: the classic Telegraph model [30] and the special case where transcription becomes a renewal process [37].
 3.1 Special Case I: Telegraph model
-Taking $D_{0}=\begin{pmatrix}-k_{1}-\lambda&k_{1}\\ k_{2}&-k_{2}\end{pmatrix}$ and $D_{1}=\text{diag}[\lambda,0]$, our model (1) reduces to the classic Telegraph model [4], the pioneering work in the field of stochastic gene expression.
+With $D_{0}=\begin{pmatrix}-k_{1}-\lambda&k_{1}\\ k_{2}&-k_{2}\end{pmatrix}$ and $D_{1}=\text{diag}[\lambda,0]$, our model (1) reduces to the classic Telegraph model, the pioneering mathematical formulation in the field of stochastic gene expression.
 It is clear that $\bm{\pi}=\left[k_{2}/(k_{1}+k_{2}),k_{1}/(k_{1}+k_{2})\right]^{T}$ and
@@
 (12) is consistent with previous result [13, 30].
-A detailed calculation can be found in the Supplementary Material [21].
+A detailed calculation can be found in Appendix F.
 Although formulation of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ in (12) seems formally compact, the confluent hypergeometric function is, in essence, an abbreviation of an infinite series.
 Evaluating the value of confluent hypergeometric function at given point possesses approximately the same computational complexity as one calculation in (7), and it involves truncation error as well.
-From this standpoint, (12) holds limited advantage, both mathematically and computationally, over reconstructing $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ from binomial moments according to (7).
+From this standpoint, (12) holds limited advantages, both mathematically and computationally, over reconstructing $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ from binomial moments according to (7).
 Likewise, this observation holds for other known results involving generalized hypergeometric function [4, 13, 16, 37].
@@
 Transcription in (1) becomes a renewal process if and only if the following renewal condition [37] holds:
-$\text{if}\;\;\sum_{k=1}^{N}b_{i,k}\neq 0,\;\;\text{then}\;\;\frac{b_{i,k}}{\sum_{k=1}^{N}b_{i,k}}\;\;\text{is independent of}\;\;i.$ (13)
-Assuming $D_{1}\neq\bm{0}_{N\times N}$, the renewal condition (13) is equivalent to $\text{rank}\,D_{1}=1$.
-Under the renewal condition, our model (1) can be interpreted otherwise as a queueing system $GI/M/\infty$ with probability density function of inter-arrival time [17]
+$\text{When}\;\;\sum_{k=1}^{N}b_{i,k}\neq 0,\;\;\frac{b_{i,k}}{\sum_{k=1}^{N}b_{i,k}}\;\;\text{is independent of}\;\;i.$ (13)
+Assuming $D_{1}\neq\bm{0}_{N\times N}$, renewal condition (13) is equivalent to $\text{rank}\,D_{1}=1$.
+Under renewal condition, our model (1) can be interpreted otherwise as a queueing system $GI/M/\infty$ with probability density function of inter-arrival time [37]
 $f(t)=\bm{\kappa}^{T}\mathrm{e}^{D_{0}t}D_{1}\bm{1},$ (14)
-where $\bm{\kappa}=(\kappa_{j})^{T}=(\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}})^{T}$ as long as $\sum_{j=1}^{N}b_{i,j}\neq 0$.
+where $\bm{\kappa}=(\kappa_{j})^{T}:=(\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}})^{T}$ as long as $\sum_{j=1}^{N}b_{i,j}\neq 0$.
 Queueing system with arrivals following a renewal process has been thoroughly studied [22].
@@
 $B_{m}=\frac{1}{\alpha m\delta}\prod_{i=1}^{m-1}\frac{\mathcal{L}[f](i\delta)}{1-\mathcal{L}[f](i\delta)},\quad m\geq 1,$ (15)
-[…]
+where $\mathcal{L}[f](s):=\int_{0}^{\infty}\mathrm{e}^{-st}f(t)\mathrm{d}t$ is the Laplace transform of $f(t)$ and $\alpha:=\int_{0}^{\infty}tf(t)\mathrm{d}t$ is the expectation of inter-arrival time.
 Assuming nonsingularity of $D_{0}$, binomial moments given in (15) can be reduced to our result given in (5).
-The details can be found in the Supplementary Material [21].
+The details can be found in Appendix G.
 We note that, under the assumption $D$ is irreducible and $\text{rank}\,D_{1}=1$, $D_{0}$ is nonsingular, which is equivalent to convergence of $\alpha$.
-A proof is given in the Supplementary Material [21], which can be seen as a generalization of Taussky theorem [23].
+A proof is presented in Appendix H, which can be seen as an analogue to Taussky theorem [15].
 3.3 Computational Performance
@@
 As dimension of $D_{0}$ or $D_{1}$ increases, number of reaction pathways escalates, making the model more complicated.
-In the above illustration, computation time pertains to an exemplar where $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$.
-In the left figure, three different approaches are employed to models with dimension $i\,(3\leq i\leq 8)$; in the right one, two approaches are employed to model with dimension $5j\,(1\leq j\leq 10)$.
+In the above illustrations, computation time pertains to exemplars where $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$.
+In the left panel, three different approaches are employed to models with dimension $k\,(3\leq k\leq 8)$; in the right one, two approaches are employed to models with dimension $5l\,(1\leq l\leq 10)$.
 Curve of our result is computation time of the first $200$ binomial moments based on (5).
@@
 Time is evaluated for three replicate computations and averaged to mitigate randomness.
-Our result for (1) is performed mainly in the framework of matrix computation, which is one of the most efficient and well-developed area in scientific computation.
-As we illustrate in Figure 3, our result stands out among several existing approaches in computational efficiency, where we compute the first $200$ binomial moments based on (5) and (15), respectively, and implement SSA.
+Our result (5) is performed mainly in the framework of matrix computation, which is one of the most efficient and well-developed area in scientific computation.
+As we illustrate in Figure 3, our result stands out in computational efficiency among several existing approaches, where we compute the first $200$ binomial moments based on (5) and (15), respectively, and implement SSA.
 Separation of time scales can be distinctly observed with respect to our result, the renewal process approach and SSA.
-In particular, our result remains computationally feasible for high-dimensional model while other approaches certainly fail.
-Moreover, our result also admits explicit error estimation, which is, as far as we know, unattainable through other existing methods.
+In particular, our result remains computationally feasible for high-dimensional model while other approaches may fail.
+Moreover, our result also admits explicit error estimation (9), which is, as far as we know, unattainable through other existing methods.
 Additionally, accuracy, stability and further improvement on efficiency can be pursued by exploiting the structure of matrix in (5), especially $k\delta\bm{I}_{N}-D$.
@@
 The second is that SSA relies heavily on empirical insights.
-Number of simulations and truncation time point to generate a stationary distribution are both empirically determined, hence the implementation requires repeated trials.
-The renewal process approach also has several shortcomings.
-Primarily, Laplace transform of inter-arrival time probability density distribution can be mathematically challenging to obtain given specific gene expression model.
-Inter-arrival time distribution also conceals physical interpretation of gene-state architecture while parameters in (1) have clear biophysical meanings.
-Additionally, results based on renewal process is computational infeasible excluding a few special cases.
-In general, computational complexity become exceedingly high if high-order binomial moment is involved.
-4 Mathematical Description of Noise in Gene Expression
+To generate a stationary distribution, number of simulations and truncation time point are both empirically determined, hence the implementation requires repeated trials.
+The renewal process approach also has shortcomings.
+Primarily, Laplace transform of inter-arrival time probability density function can be mathematically challenging to obtain given specific gene expression model.
+Inter-arrival time probability density function also conceals physical interpretation of gene-state architecture while parameters in (1) have clear physical meanings.
+Additionally, results based on renewal process are computational infeasible excluding a few special cases, because, in general, computational complexity becomes exceedingly high if high-order binomial moments are involved.
+4 Fluctuation of mRNA Copy Number in Transcription
 Noise in gene expression within a supposedly identical ensemble arises from multiple sources [26, 32].
-Quantitative understanding of generation and propagation of noise is one of the primary objectives of studying stochastic gene expression, and noise in transcription can predominantly contribute to protein-level fluctuation [29].
+Quantitative understanding of generation and propagation of noise is one of the primary objectives of modeling stochastic gene expression, and noise in transcription can predominantly contribute to protein-level fluctuation [28].
 Experimentally, noise can be categorized into intrinsic and extrinsic noise.
@@
 Deterministic description fails since number of molecules involved in gene expression can be relatively low.
-Our model (1) provides a general framework to investigate origins and consequences of intrinsic noise in gene expression.
+Our model (1) provides a general framework of investigating origins and consequences of intrinsic noise in transcription.
 Based on our result (5), Fano factor, the ratio of variance to mean of mRNA copy number, is given by
-$FF=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}}=\frac{\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{\bm{\pi}^{T}D_{1}\bm{1}}+1-\frac{\bm{\pi}^{T}D_{1}\bm{1}}{\delta}.$ (16)
+$FF:=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}}=\frac{\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{\bm{\pi}^{T}D_{1}\bm{1}}+1-\frac{\bm{\pi}^{T}D_{1}\bm{1}}{\delta}.$ (16)
 Intrinsic noise, defined as the ratio of variance to mean square, is
-$\eta^{2}=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}^{2}}=\frac{\delta\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{(\bm{\pi}^{T}D_{1}\bm{1})^{2}}+\frac{\delta}{\bm{\pi}^{T}D_{1}\bm{1}}-1.$ (17)
+$\eta^{2}:=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}^{2}}=\frac{\delta\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{(\bm{\pi}^{T}D_{1}\bm{1})^{2}}+\frac{\delta}{\bm{\pi}^{T}D_{1}\bm{1}}-1.$ (17)
 For specific gene-state structure where $(\delta\bm{I}_{N}-D)^{-1}$ has explicit expression, further reduction of Fano factor and intrinsic noise may be performed.
 Consistency of (16) and (17) with previous results can be verified [4, 37].
+Derivation of (16) and (17) is in Appendix I.
+As illustrated in Figure 4, intrinsic noise converges to zero as model size increases, implying that kinetics of mRNA molecules tends towards deterministic description in these exemplars.
+Additionally, Fano factor is greater than $1$ except for the trivial $n=1$ case according to Figure 4, meaning that distribution of mRNA copy number is super-Poissonian in these exemplars.
+Figure 4: Variation of Intrinsic Noise and Fano Factor: Intrinsic noise and Fano factor are computationally evaluated for models of order $n$($1\leq n\leq 100$), with parameters $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$.
+In the left panel, intrinsic noise $\eta^{2}$ is computed based on (17) and $-\ln(\eta^{2})$ is taken for plotting.
+In the right panel, Fano factor is computed based on (16).
 Extrinsic noise refers to fluctuation caused by environmental differences among cells, and can be further classified as local and global extrinsic noise [32].
@@
 One insightful approach is to utilize the parametric nature of the model [10].
-Parameters, which are $D_{0},D_{1}$ and $\delta$ in our case, are assumed to follow some probability distribution with density function $\rho(\bm{\theta})$, representing effects of extrinsic fluctuation.
-Since binomial moments have concise expression given by (5), posing the perturbation on binomial moments and then reconstruct probability distribution is more tractable than posing perturbation directly on probability distribution as in [19].
+Parameters, which are $D_{0},D_{1}$ and $\delta$ in our case, are assumed to follow given probability distribution with density function $\rho(\bm{\theta})$, representing effects of extrinsic perturbation.
+Since binomial moments have concise expression given by (5), posing the perturbation on binomial moments and then reconstruct probability distribution is more tractable than posing perturbation directly on probability density function as in Ref.
+[10].
 Thereby, the expected binomial moments of mRNA molecules subject to extrinsic noise can be formulated as
-[…]
+$\langle B_{m}\rangle_{\rho}=\int_{\Omega}B_{m}(\bm{\theta})\rho(\bm{\theta})\mathrm{d}\bm{\theta},$ (18)
 where $\bm{\theta}$ abbreviates the vector of $12N^{2}-N+$ 1 independent parameters in our model (1) and $\Omega$ denotes the parameter space $[0,\infty]^{2N^{2}-N+1}$.
-Substituting (18) into (7), perturbed probability distribution can be reconstructed.
+Substituting (18) into (7), perturbed probability density function can be reconstructed.
 For specific $\rho(\bm{\theta})$ and simplified model, (18) may be reduced to closed form.
+In general, (18) can also be numerically evaluated by Monte Carlo integration since computation of $B_{m}$ is relatively fast.
 This framework can be used to apprehend the joint effects of intrinsic and extrinsic noise, which has been performed for some special cases of our model (1) [10, 13].
-Heavy-Tailed Law [19] asserts that intrinsic noise alone cannot generate mRNA distributions that is heavy tailed, namely, the moment generating function [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)$ becomes infinity for $t>0$.
+Heavy-Tailed Law asserts that intrinsic noise alone cannot generate mRNA distributions that is heavy tailed, namely, the moment generating function [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)$ becomes infinity for any $t>0$.
 Heavy-Tailed Law has been proposed intuitively for Leaky Telegraph model [10] and a generalized model [13].
-In this Letter, we provide a mathematically rigorous proof of Heavy-Tailed Law in our model (1), generalizing previous results.
-More preciously, we prove the following inequality based on (8).
+Here we provide a mathematically rigorous proof of Heavy-Tailed Law in our model (1), generalizing previous results.
+Denote Poisson distribution with parameter $\Lambda$ by $\text{Pois}(\Lambda)$.
+We prove the following inequality based on (8).
 [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)\leq\mathrm{e}^{2\lVert D_{1}\rVert_{\infty}/\delta}\times\mathcal{M}\left[\text{Pois}(\frac{\lVert D_{1}\rVert_{\infty}}{\delta})\right](t).$ (19)
 For Poisson distribution, moment generating function is finite for any $t>0$, therefore so is $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ according to (19).
-Therefore $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is never heavy tailed.
 It follows that experimentally observed heavy-tailed distribution of mRNA copy number implies collaborative effect of intrinsic and extrinsic noise [10].
-The mathematical proof of (19) is in Supplementary Material [21].
+The mathematical proof of (19) is in Appendix J.
 5 Discussion
@@
 Numerous mathematical theories have been developed to provide quantitative insights into specific models.
-In this Letter, we study a general kinematic model of transcription.
-The exact time-dependent solution to the chemical master equation is constructed.
-A computationally feasible approach based on binomial moments is presented and thoroughly analyzed.
-We examine two special cases and perform SSA to verify our results.
-Collaborative effect of intrinsic and extrinsic noise are discussed.
-A mathematically rigorous proof of generalized Heavy-Tailed Law is also provided.
-Compared with other existing methods, our approach features generality, exceptional computational efficiency as well as explicit error estimation.
-Results presented in this Letter boost further analysis of origin and variation of noise.
+This article establishes a general framework for studying stochastic transcription kinetics and related problems.
+While pursuing the utmost theoretical generality, we construct the exact time-dependent solution to CME and present a computationally feasible approach based on binomial moments.
+We unify existing studies on transcription kinetics and discuss the combined effects of intrinsic and extrinsic noise.
+Additionally, a mathematically rigorous proof of generalized Heavy-Tailed Law is provided.
+Compared with other existing methods, our approach features generality, exceptional computational efficiency as well as explicit truncation error estimation.
+Results presented in this article boost further analysis of origin and variation of noise.
 Dependence of noise and Fano factor on different parameters should be further investigated.
@@
 When modeling specific type of gene, physical background should be exploited to constrain model complexity appropriately.
-Statistical inference can further be employed by fitting our model to experimental data, and finally reveal the underlying mechanisms of transcription.
+Statistical inference can further be employed by fitting our model to experimental data, eventually revealing the underlying mechanisms of transcription.
 Acknowledgments
@@
 L.
-is supported by FDUROP (Fudan Undergraduate Research Opportunities Program).
+is supported by FDUROP (Fudan Undergraduate Research Opportunities Program)(24260).
 Y.
@@
 is supported by National Key R&\&&D Program of China (2024YFA1012401), the Science and Technology Commission of Shanghai Municipality (23JC1400501), and Natural Science Foundation of China (12241103).
+Author Declaration
+Conflict of Interest
+The authors have no conflicts to disclose.
+Author Contributions
+Yuntao Lu: Conceptualization (equal); Formal analysis (lead); Software (lead); Visualization (lead); Writing - original draft (lead); Writing - review & editing (equal).
+Yunxin Zhang: Conceptualization (equal); Project administration (lead); Supervision (lead); Writing - review & editing (equal).
+Data Availability Statement
+The data that supports the findings of this study are available within the article.